OpenStudy (anonymous):

Use the following equation: y = –3(x + 8)^2 + 2 The coordinates of the vertex written as (h, k) would be? a. (–8, –3) b. (8, 3) c. (–8, 2) d. (8, –2)

OpenStudy (anonymous):

@ankit042 :)

OpenStudy (antonio_xx2):

y=a(x-h)^2=k vertex (h,k) y=-3(x+8)^2=2 figure it out;)